# SPARC: Prediction-Based Safe Control for Coupled Controllable and Uncontrollable Agents with Conformal Predictions

### v1

@@ IV-D Safe Control using Control Barrier Functions
After bounding the state estimation error in Proposition 1, we calculated the conformal prediction region bound $\bar{R}$ for Eq. (6) to design a uncertainty-aware controller via the technique of control barrier functions.
@@
Therefore, Problem 1 can be formulated as a safe control problem within each prediction update period by considering the estimated region of the uncontrollable agent as an obstacle simply by define
$\mathcal{C}=\{x\in\mathbb{R}^{n}\mid\forall Y\in\mathbb{R}^{n_{Y}}:(x,Y)\in\mathcal{S}\}$ (9)
as the safe set for the controlled system such that the overall system remain in safe set $\mathcal{S}$. Therefore, it suffices to find a CBF $\hat{h}(x)$ such that
$\hat{C}=\{x\in\mathbb{R}^{n}\mid\hat{h}(x)\geq 0\}\subseteq\mathcal{C}.$ (10)
In the literature, there are many existing approach for finding such CBF $\hat{h}(x)$, e.g., by sum-of-square programming [26] or learning-based approach [27, 28].
Now, given the synthesized CBF $\hat{h}(x)$, the set of admissible control inputs is defined by
$K_{CBF}:=\{u\in\mathcal{U}\mid L_{f}\hat{h}(x)+L_{g}\hat{h}(x)u\geq-\beta(\hat{h}(x))\}$ (11)
The following result shows that, by following the control strategy in Eq. (11), one can ensure safety within each period $\Delta$ with confidence $1-\alpha$.
@@
Consider the control law $u(t)$ from $K_{CBF}$ in Eq. (11),the system composed by controllable and uncontrollable variables Eq. (1), Eq. (2) satisfies Eq. (8) where $\mathcal{L}_{L_{f}h},\mathcal{L}_{\beta\circ h}$, and $\mathcal{L}_{L_{g}h}$ are the Lipschitz constants of the functions $L_{f}h,\beta\circ h$ and $L_{g}h$, respectively, we have
@@
When $Y_{t+1}\in\hat{\mathcal{O}}_{Y_{t+1}}(\alpha)$ (with probability higher than $1-\alpha$), the“obstacle”(i.e., the safe region which includes the actual position of the uncontrollable agent inside) is accurately estimated. Consequently, the control law $u(t)$ derived from $K_{CBF}$ for $\hat{\mathcal{C}}$ is effective in ensuring the safety of the entire system, meaning that $(X_{t+1},Y_{t+1})$ remains within the safe set $\mathcal{S}$, as defined in Problem 1. ∎
@@ V Simulation Results
In this section, we illustrate and evaluate the proposed approach through simulations of a case study involving autonomous driving with uncontrollable pedestrians. This case study demonstrates the effectiveness of our framework in managing the interactions between the autonomous vehicle and pedestrians, ensuring safety despite the uncertainty in pedestrian behavior.
Figure 3: The pedestrian’s walking logic is modeled as an unknown stochastic distribution affected by the state of the vehicle, i.e.,$\Delta Y\sim\mathcal{D}(X)$. The pedestrian starts at a random position on one side of the crosswalk while the vehicle begins at the midpoint of the starting line. Red circles indicate collision volume.
Figure 4: Nonconformity scores histogram for pedestrian position displacement $(\Delta x,\Delta y)$ on $\mathcal{D}_{cal}$
V-A Scenario Description and System Dynamic
@@
Vehicle Dynamics: The vehicle’s starting speed is sampled from a uniform distribution within the range $[0,15]$ m/s, which mimics typical vehicle speeds observed in urban environments (up to the speed limit of 15 m/s). The vehicle’s position is updated based on its velocity.
Pedestrian Behavior: The pedestrian’s movement is modeled as a unknown distribution of the vehicle’s state-space vector and Gaussian noise i.e.
$\Delta{Y}=\phi(X)+\mathcal{N}({0,\sigma})\sim\mathcal{D}(X)$ (13)
@@
Interaction and Collision Avoidance: The pedestrian’s trajectory is evaluated in real-time as the vehicle approaches. We track the pedestrian’s crossing behavior when the pedestrian is on the crosswalk, where the vehicle’s speed directly impacts the pedestrian’s decision to speed up, slow down, or stop. The simulation continuously monitors the distance between the vehicle and the pedestrian, and the safe controller(SPARC) is applied to ensure safe control synthesis in avoiding collisions.
Trajectory Predictor: We trained a feedforward neural network to predict the future trajectory of the pedestrian. The input is the history state measurement of the vehicle’s and pedestrian’s state. The training dataset $D_{\text{train }}$ contains $1\times 10^{6}$ data points, and the calibration dataset $D_{\text{cal }}$ for conformal prediction contains $1\times 10^{5}$ data points.
Barrier Functions: For vehicle-pedestrian collision avoidance, we may define the CBF as follows:
$h(x)=\min_{Y\in\partial\hat{\mathcal{O}}_{Y}}||x-Y||-d_{\text{safe}}$ (14)
where:
  • $x\in\mathbb{R}^{n}$ is the position of the vehicle
  • $\partial\hat{\mathcal{O}}_{Y}\subseteq\mathbb{R}^{n_{Y}}$ is the predicted safe region boundary of the uncontrollable pedestrian
  • $d_{\text{safe}}$ is the minimum safe distance to be maintained
[…]
This CBF in a sense of probability ensures a safe distance is maintained between the vehicle and the pedestrian. The system is in a safe state, avoiding collision, when $h(x)\geq 0$. The controller design should satisfy the inequality in Eq. (8) to guarantee system safety. Set $d_{\text{safe}}$ = 1.0 m, the collision volume is shown as the red circle in Fig. 3.
@@
    Uncontrollable Prediction Robust CBF: Let $\alpha$ be the failure tolerance, randomly pick $\alpha=0.85,0.75,0.5$, the results are shown in Table I.
@@
    Vanilla CBF: We set $\alpha=0\implies\bar{R}(\alpha)=[0,0,\ldots,0]$, which essentially reduces to the vanilla CBF [8]. The barrier function is directly from neural network output rather than safe region, so this CBF cannot provide any safety guarantee.
V-B Result Analysis
Figure 5: The result of the conformal prediction in the case study
Trajectory Predictors and Conformal Prediction: Take failure probability tolerance $\alpha=0.85$ as an example, Fig. 4 illustrates a calculation of the nonconformity scores used in our approach. The scores $R_{x}^{0.85}=0.75,R_{y}^{0.85}=2.86$ represent the deviation between the predicted pedestrian position change $\Delta\hat{Y}=(\Delta x,\Delta y)$ and the ground truth, providing a measure of prediction uncertainty , which is used to generate safe control decisions.
Fig. 5 shows the network successfully learns the underlying random distribution governing pedestrian behavior in relation to the vehicle’s state. In our simulation framework, for the sake of simplicity, the upper limit of all nonconformity scores is employed as a proxy for the overall safety margin, and it is also feasible to utilize a grid-based distribution of nonconformity scores to devise a more aggressive control strategy.
Safe Control Barrier Function Controller: The vehicle driven by our SPARC is expected to avoid collisions with pedestrians. We set the failure probability tolerance $\alpha=0.15,0.25,0.5$, and desire $\text{Prob}[(X,Y)\in\mathcal{S}]\geq 1-\alpha$, which means $\text{Prob}[\text{vehicle and pedestrian collide}]\leq\alpha$. We performed $1\times 10^{5}$ experiments, each time step is set to 100, the collision probability are collected in Table I. In 11.20% of the trials(1120 experiments), the car encounter a pedestrian, and we care about whether our controller can successfully prevent collision in these experiment, so we divide absolute collision frequency by 11.20% to get relative collision frequency. The collision rate of SPARC are $14.46\%(<15\%),17.50\%(<25\%),37.59\%(<50\%)$, which are all under their respective failure probability tolerance and significantly lower than vanilla CBF case $(96.43\%)$, so our Proposition 2. holds in practice.
@@ VI Conclusion
This work presents a new framework that integrates trajectory prediction with conformal prediction to enhance safety in multi-agent systems involving coupled controllable and uncontrollable agents. By incorporating probabilistic guarantees into the control barrier function framework, we ensure safe interactions between the controlled system and uncontrollable agents with high confidence. Compared to previous approaches, our method accounts for the coupling between the controlled system and the uncontrollable agent by sampling and evaluating over an augmented state space, making it more applicable to real-world scenarios such as autonomous driving and human-robot interaction. However, one limitation of our approach is the assumption that precise state information for both the controlled system and uncontrollable agent is fully accessible. In many practical applications, such state information must be obtained through perception devices like LiDAR, which can introduce estimation errors. In the future, we plan to extend our framework to perception-based settings, addressing the additional challenge of state estimation uncertainty.

### v3

@@ IV-D Safe Control using Control Barrier Functions
After we quantified the uncertainty with the calculated prediction region, we design an uncertainty-aware safety filter with CBF.
@@
At each time step, the state of the uncontrollable agent is treated as a fixed parameter, reducing $h_{xy}(x,y)$ to $h(x)$. Therefore, Problem 1 can be formulated as a safe control problem within each prediction update period by considering the estimated region of the uncontrollable agent as an obstacle simply by define
$\mathcal{C}=\{x\in\mathbb{R}^{n}\mid\forall y\in\mathbb{R}^{m}:(x,y)\in\mathcal{S}\}$ (12)
as the safe set for the controlled system such that the overall system remain in safe set $\mathcal{S}$. Therefore, it suffices to find a CBF $h(x)$ such that
$C=\{x\in\mathbb{R}^{n}\mid{h}(x)\geq 0\}\subseteq\mathcal{C}.$ (13)
In the literature, there are many existing approach for finding such CBF $h(x)$, e.g., by sum-of-square programming [26] or learning-based approach [27, 28].
Now, given the synthesized CBF $h(x)$, the set of admissible control inputs is defined by
$K_{CBF}:=\{u\in\mathcal{U}\mid L_{f}h(x)+L_{g}h(x)u\geq-\beta(h(x))\}$ (14)
The safe input can be any control law $u(t)$ from $K_{CBF}$ in Eq. (14), e.g., one computed by solving the quadratic optimization problem (3).
@@
Consider the control law $u(t)$ from $K_{CBF}$ in Eq. (14),the system composed by controllable and uncontrollable variables Eq. (1), Eq. (4) satisfies Eq. (11), respectively, we have
@@
With guarantee that $\operatorname{Prob}[Y_{t+1}\in\hat{\mathcal{O}}_{Y_{t+1}}(\alpha)]\geq 1-\alpha$, the control law $u(t)$ derived from $K_{CBF}$ for $\hat{\mathcal{C}}$ is effective in ensuring the safety of the entire system, meaning that $(X_{t+1},Y_{t+1})$ remains within the safe set $\mathcal{S}$, as defined in Problem 1. ∎
@@ V Simulation Results
In this section, we illustrate and evaluate the proposed approach through simulations of a case study involving autonomous driving with uncontrollable pedestrians. This case study demonstrates the effectiveness of our framework in modelling the interactions between the autonomous vehicle and pedestrians, providing probabilistic safety guarantees under the uncertainty of pedestrian behavior.
Figure 3: The pedestrian’s speed follows an unknown stochastic distribution affected by the state of the vehicle, i.e.,$\Delta Y\sim\mathcal{D}(X,Y)$. The pedestrian starts at a random position on one side of the crosswalk while the vehicle begins at the midpoint of the starting line. Red circles indicate collision volume.
Figure 4: Nonconformity scores histogram for pedestrian position displacement $(\Delta{Y_{par}},\Delta{Y_{perp}})$ on $D_{cal}$
V-A Scenario Description and System Dynamics
@@
Vehicle Dynamics: For the sake of simplicity, we assume the vehicle moves straight through the crosswalk, setting the dimension $n=1$ with dynamic model $\dot{x}=u$. This choice aligns with real-world constraints, as vehicles rarely make left or right turns while crossing a zebra line. Nevertheless, our approach naturally extends to two or higher dimensions. The vehicle’s speed range is $[0,15]$m/s, reflecting typical vehicle speeds observed in urban environments (up to the speed limit of 15 m·s-1/54 km·h-1/33 mph). The vehicle’s position is updated based on its velocity.
Pedestrian Behavior: The pedestrian’s movement is modeled as a stochastic process governed by a function of the vehicle’s state-space vector and Gaussian noise. Mathematically, this can be expressed as:
$\Delta{Y}\sim\mathcal{N}_{2}(\phi(X,Y),\Sigma),$ (16)
where $\Delta Y=(\Delta{Y_{par}},\Delta{Y_{perp}})\in\mathbb{R}^{2}$ follows a Gaussian distribution whose mean value is a function of both the state of the pedestrian and the controlled vehicle. Here the notation $\Delta{Y_{par}}$ represents the velocity component parallel to the road direction, while $\Delta{Y_{perp}}$ denotes perpendicular component. $\Sigma$ is the standard deviation $\Sigma=\operatorname{diag}(\sigma^{2}_{x},\sigma^{2}_{y})=\operatorname{diag}((\text{0.5m/s})^{2},(\text{2m/s})^{2})$. To illustrate, as the vehicle approaches at a higher speed, the pedestrian slows down $\Delta{Y_{perp}}$, exhibiting more cautious crossing behavior. Simultaneously, $\Delta{Y_{par}}$ increases in the direction of moving away from the vehicle as an evasive response. Additionally, $\Delta{Y_{par}}$ is influenced by $Y$. When approaching the left/right edge of the crosswalk area, the pedestrian tends to adjust the position toward the center of the crosswalk. And as pedestrian movement is inherently unpredictable, this modelling introduces additional stochasticity to account for hesitation and spontaneous decision-making.
Under this simulation setup, pedestrians exhibit a perturbed yet generally arced crossing trajectory around the vehicle. This behavior aligns with the natural tendency of pedestrians to adjust their paths dynamically while crossing.
V-B Model Implementation
Interaction and Collision Avoidance: The pedestrian’s trajectory is evaluated in real-time as the vehicle approaches. We track the pedestrian’s crossing behavior which is influenced by the vehicle. The simulation continuously monitors the distance between the vehicle and the pedestrian, and the safe controller (SPARC) is applied to ensure safe control synthesis in avoiding collisions.
Trajectory Predictor: We trained a feedforward neural network to predict the future trajectory of the pedestrian. The dataset is generated by simulating a vehicle moving at a constant speed while tracking both the vehicle and pedestrian trajectories. The vehicle’s speed is uniformly sampled from the range $[0,15]$m/s. The collected data is structured following the format described in the IV. B section. The input is the history state measurement of the vehicle’s and pedestrian’s state. The training dataset $D_{\text{train }}$ contains $1\times 10^{6}$ data, and the calibration dataset $D_{\text{cal}}$ for conformal prediction contains $1\times 10^{5}$ data. The model consists of three fully connected layers with ReLU activation. The input layer maps features to 64 hidden units, followed by a second layer of 64 units, and a final output layer predicting pedestrian speed $(\Delta{Y_{par}},\Delta{Y_{perp}})\in\mathbb{R}^{2}$. The model use the Adam optimizer (learning rate 0.001) and MSE loss. A StepLR scheduler halves the learning rate every 10 epochs over 50 epochs.
Control Barrier Functions: For vehicle-pedestrian collision avoidance, we choose the following CBF:
$h(x)=\min_{y\in\hat{\mathcal{O}}_{y}}||x-y||-d_{\text{safe}}$ (17)
where:
  • $x\in\mathbb{R}^{n}$ is the position of the vehicle,
  • $\hat{\mathcal{O}}_{y}\subseteq\mathbb{R}^{m}$ is the predicted conformal region of the uncontrollable pedestrian, and we set $\hat{\mathcal{O}}_{y}(\alpha)=[\hat{y}-\bar{R}(\alpha),\hat{y}+\bar{R}(\alpha)]\subseteq\mathbb{R}^{2}$ as a weighted $\ell_{\infty}$-norm ball.
  • $d_{\text{safe}}$ is the minimum safe distance to be maintained,
@@ V-B Model Implementation
With $d_{\text{safe}}$ = 1.0 m, the collision volume is shown as the red circle in Fig. 3.
@@
    SPARC: Let $\alpha$ be the failure tolerance, randomly pick $\alpha=0.85,0.75,0.5$, the results are shown in Table I.
@@
    Vanilla CBF: We set $\alpha=1$, which means that $\bar{R}(\alpha)=[0,0]$. This essentially reduces to the vanilla CBF [8]. This method uses the direct output from the neural network output than the calibrated conformal region, providing no safety guarantee and failing to account for dynamic stochasticity.
  3. 3.
    Random: In this experiment, the vehicle’s initial velocity is uniformly sampled, and it follows the reference control $u_{r}$ during its motion. The proportion of trials resulting in a collision with the pedestrian is recorded.
@@
Figure 5: The prediction output with conformal prediction region. The plot also includes $\phi(\text{car speed})$ with all other parameters fixed as a noise-free reference ground truth.
V-C Result Analysis
Trajectory Predictors and Conformal Prediction: Take failure probability tolerance $\alpha=0.85$ as an example, Fig. 4 illustrates a calculation of the nonconformity scores used in our approach. The scores $R_{Y_{par}}^{0.85}=0.75,R_{Y_{perp}}^{0.85}=2.86$ represent the deviation between the predicted pedestrian position change $\Delta\hat{Y}$ and the ground truth, providing a measure of prediction uncertainty , which is used to generate safe control decisions.
Fig. 5 shows the how the network learns the underlying random distribution governing pedestrian behavior in relation to the vehicle’s state. Notably, our method does not rely on the precise accuracy, overall performance, or scene comprehension capabilities of the model, making it robust to model imperfections and uncertainties. In our simulation framework, for the sake of simplicity, the upper limit of all nonconformity scores is employed as a proxy for the overall safety margin, and it is also feasible to utilize a grid-based distribution of nonconformity scores to devise a more aggressive control strategy.
Safe Control Barrier Function Controller: The vehicle controlled by our SPARC framework is expected to avoid collisions with pedestrians. We set the failure probability tolerance $\alpha=0.15,0.25,0.5$, and desire $\text{Prob}[(X,Y)\in\mathcal{S}]\geq 1-\alpha$, which means $\text{Prob}[\text{vehicle and pedestrian collide}]\leq\alpha$. We performed $1\times 10^{5}$ experiments, each time step is set to 100, the collision probability are collected in Table I. In 11.20% of the trials (1120 experiments), the car crash on the pedestrian, and we care about whether our controller can successfully prevent collision in these experiment, so we divide absolute collision frequency by 11.20% to get relative collision frequency. Due to the relatively high ratio of the standard deviation of the added stochastic component to the pedestrian’s average speed, the vanilla CBF, which relies solely on predicted pedestrian positions for avoidance, struggles to account for this uncertainty and consequently suffers from a high failure rate$(96.43\%)$. The collision rate of SPARC are $14.46\%(<15\%),17.50\%(<25\%),37.59\%(<50\%)$, which are all under their respective failure probability tolerance and significantly lower than vanilla CBF case $(96.43\%)$, so our Proposition 2. holds in practice.
@@ VI Conclusion
This work presents a new framework that integrates trajectory prediction with conformal prediction to enhance safety in multi-agent systems involving coupled controllable and uncontrollable agents. By incorporating probabilistic guarantees into the control barrier function framework, we ensure safe interactions between the controlled system and uncontrollable agents with high confidence. Compared to previous approaches, our method accounts for the coupling between the controlled system and the uncontrollable agent by sampling and evaluating over an augmented state space, making it more applicable to real-world scenarios such as autonomous driving and human-robot interaction.
There are several future directions for this work that we plan to explore. The main objective of our work is to present a generic framework applicable to a broad class of control synthesis problems involving coupled uncontrollable agents, rather than being tailored to specific applications. In this work, for illustrative purposes, we use simplified vehicle and pedestrian models to demonstrate our approach. Regarding the complex nature of vehicle-pedestrian interactions, various datasets and prediction models have been developed to capture realistic behaviors. Notable datasets such as ETH & UCY [29] and JAAD [30] provide benchmarks for pedestrian trajectory modeling, while methods like Social-LSTM [31] and Trajectron++ [32] offer deep learning-based approaches for trajectory prediction. The TrajNet++ benchmark [33] further standardizes evaluation in this domain. In the future, we plan to focus specifically on autonomous driving by leveraging these real-world datasets and customized prediction methods to enhance the accuracy and applicability of our framework. Additionally, one limitation of our current approach is the assumption that precise state information for both the controlled system and the uncontrollable agent is fully accessible. In many practical applications, such state information must be obtained through perception devices like LiDAR, which can introduce estimation errors. To address this, we plan to extend our framework to perception-based settings, incorporating techniques to handle state estimation uncertainty and further improve the robustness of our approach in real-world scenarios.